Detecting multiparton interactions in minimum-bias events at ALICE

Raffaele Grosso^{1}^{†}, Jan Fiete Grosse-Oetringhaus^{2} for the ALICE collaboration 1INFN, Sezione di Torino, Italy 2CERN – European Organization for Nuclear Research

Abstract

The observed long tail of high-multiplicity events has questioned the current modelizations for the charged-particle multiplicity distribution. It has been interpreted as an indirect observation of multiparton interactions becoming increasingly important at higher collision energies. The ALICE detector will measure the frequency of very highmultiplicity events. The performance for measuring the charged-particle multiplicity distribution in ALICE is presented.

1 Introduction

Being at LHC the heavy-ion dedicated experiment, ALICE – A Large Ion Collider Experiment [1] – has some unique capabilities, complementary to those of the dedicated p-p experiments. Its 18 detector systems have been designed to provide high-momentum resolution as well as excellent Particle Identification (PID) over a broad momentum range (in particular with very low p_{T} -cutoff) and up to the highest multiplicities predicted for LHC.

Besides running with Pb ions, the physics programme includes collisions with lighter ions, lower energy running and dedicated proton-nucleus runs. ALICE will also take data with proton beams at varying energies, up to the top LHC energy, to collect reference data for the heavy-ion programme and to address several QCD topics for which ALICE is complementary to the other LHC detectors.

The charged-particle multiplicity distribution is among the measurements which are expected to shed light on the dynamics of multiparton interactions. We recall here the results of a study for evaluating the performances of measuring the charged-particle multiplicity distribution with the ALICE detector.

The frequency of non-jet events with very high multiplicity observed by CDF [2] has questioned the models for multiparticle production. Multiparton scattering increases the number of soft particles both in minimum-bias events and in the underlying event associated with high-p_{t} jets. It is expected that multiparton interactions are responsible for the high-multiplicity tails that break Koba-Nielsen-Olesen (KNO) [3] scaling and become significantly more important at LHC energies. The ALICE detector can make use of its very low-p_{T} cutoff (p_{T} ≈ 100 MeV) and of its high-multiplicity trigger to investigate the production of large numbers of soft particles in minimum-bias events.

†speaker (present affiliation: Università di Padova, Italy)